# Pure Lovelock Gravity regular black holes

### v1

@@ Abstract
We present a new family of regular black holes (RBH) in Pure Lovelock gravity, where the energy density is determined by the gravitational vacuum tension, which varies for each value of $n$ in each Lovelock case. A notable feature of our model is that the regular solution closely resembles the vacuum solution before reaching the event horizon. For odd $n$, the transverse geometry is spherical, with phase transitions occurring during evaporation, and the final state of this process is a remnant. For even $n$, the transverse geometry in non trivial and corresponds to a hyperboloid. In the case of $d=2n+1$ with even $n$, we find an RBH without a dS core and no inner horizon (whose presence has been recently debated in the literature due to the question of whether its presence is unstable or not), and no phase transitions. For $d>2n+1$ with even $n$, the RBH possesses both an event horizon and a cosmological horizon and no inner horizon. The existence of the cosmological horizon arises without the usual requirement of a positive cosmological constant. From both numerical and analytical analysis, we deduce that as the event horizon expands and the cosmological horizon contracts, thermodynamic equilibrium is achieved in a remnant when the two horizons coincide.
@@ I Introduction
Nowadays, it is well-accepted that black hole physics and gravity are far from being finished and several core aspects are still unclear. For instance, since the original works by Penrose and Hawking, it is well-accepted that, considering only the known quantum phenomena, gravitational collapse could lead to black hole solutions containing singularities. In this connection, it is well established that the infinite tidal forces near a black hole’s singularity can lead to the infinite stretching of an object, a phenomenon known as spaghettification. See for example the introduction of Hong et al. (2020). However, this prediction now is sometimes disregarded as the introduction of potential new quantum (gravity) effects could change the scenario drastically, avoiding the formation of singularities.
@@
Dymnikova proposes a model that encodes the gravitational information of the spherically symmetric vacuum solution in such a way that near the central singularity—where the gravitational tension and tidal forces of the vacuum black hole become infinite—the energy density in the Dymnikova model (DyM) remains finite. This premise will be crucial for the model we examine in this work. This finite density in the DyM behaves as a positive cosmological constant near the origin, resulting in a regular black hole solution with a de Sitter core. It is also worth noting that the DyM has garnered significant attention in recent years for addressing various issues in physics Paul (2023); Konoplya and Zhidenko (2024); Estrada and Muniz (2023); Alencar et al. (2023); Estrada and Aros (2019).
@@
A special case of Lovelock theories is Pure Lovelock theory (PL) Cai and Ohta (2006). As will be discussed below, this theory considers a single term in the Lagrangian. Pure Lovelock theory has drawn attention in recent years for various problems in physics Singha and Biswas (2024); Paithankar and Kolekar (2023); Shaymatov and Dadhich (2022). As we will see below, the value of the Kretschmann scalar for the vacuum solution depends on the power $n$ of the curvature. This makes it interesting to study the behavior of a non-vacuum model that encodes the gravitational information of the empty geometry in an analogous way to the Dymnikova model.
On the other hand, vacuum solutions of the motion equations in Pure Lovelock theory are doubly degenerate for even $n$, a situation that does not occur in General Relativity where $n=1$. This also raises an intriguing question about what these degenerate solutions represent in the non-vacuum case and whether they allow for transverse geometries that differ from spherical symmetry. Additionally, it is well known that General Relativity has a non-trivial vacuum solutions, where the gravitational potential of the form $-\bar{M}/r^{d-2n-1}$ does not depend radially on its denominator when $d=3$ (i.e., $d=2n+1$ where $n=1$). An interesting feature is that Pure Lovelock theory retains this property for $d=2n+1$ with $n>1$ Dadhich (2016). This property raises the question of what the physical interpretation would be in the non-vacuum case, i.e., when $d=2n+1$ in the presence of high curvature terms.
@@ IV Our regular black hole solution in Pure Lovelock gravity
    $\lim_{r\to\infty}\mu(r)_{\gamma=1}=1-\left(\frac{2(n-1)!M}{r^{d-2n-1}}\right)^{\frac{1}{n}},$ (28)
    $\lim_{r\to\infty}\mu(r)_{\gamma=-1}=-1+\left(\frac{2(n-1)!M}{r^{d-2n-1}}\right)^{\frac{1}{n}}$ (29)
@@
    $\mu(r)_{\gamma=1}\Big{|}_{r\ll a}1-\left(\frac{2M}{a^{d-1}\cdot n}\right)^{\frac{1}{n}}r^{2}.$ (30)
@@ V.1.1 $d=2n+1$
Figure 1:
@@ V.1.2 $d>2n+1$
Figure 2:
Figure 3:
@@
Figure 5:
@@ V.2.2 $d>2n+1$
Figure 7:
@@ VI.1.1 $d>2n+1$
Figure 8:
@@ VI.2.2 $d>2n+1$
Figure 9:
@@ VIII.2.1 $d=2n+1$
Figure 10:

### v3

@@ Abstract
We present a new family of regular black holes (RBH) in Pure Lovelock gravity, where the energy density is determined by the gravitational vacuum tension, which varies for each value of $n$ in each Lovelock case. Speculatively, our model may capture quantum effects through gravitational tension. In this way, a hypothetical analogy is drawn between the pair production ratio in the Schwinger effect and our energy density. A notable feature of our model is that the regular solution closely resembles the vacuum solution before reaching the event horizon. For odd $n$, the transverse geometry is spherical, with phase transitions occurring during evaporation, and the final state of this process is a remnant. For even $n$, the transverse geometry is non trivial and corresponds to a hyperboloid. In the case of $d=2n+1$ with even $n$, we find an RBH without a dS core and no inner horizon (whose presence has been recently debated in the literature due to the question of whether its presence is unstable or not), and no phase transitions. For $d>2n+1$ with even $n$, the RBH possesses both an event horizon and a cosmological horizon, also with no inner horizon present. The existence of the cosmological horizon arises without the usual requirement of a positive cosmological constant. From both numerical and analytical analysis, we deduce that as the event horizon expands and the cosmological horizon contracts, thermodynamic equilibrium is achieved in a remnant when the two horizons coincide.
@@ I Introduction
Nowadays, it is well-accepted that black hole physics and gravity are far from being finished and several core aspects are still unclear. For instance, since the original works by Penrose and Hawking, it is well-accepted that, considering only the known quantum phenomena, gravitational collapse could lead to black hole solutions containing singularities. In this connection, it is well established that the infinite tidal forces near a black hole’s singularity can lead to the infinite stretching of an object, a phenomenon known as spaghettification. See for example the introduction of Hong et al. (2020). However, this prediction now is sometimes disregarded as the introduction of potential new quantum (gravity) effects could change the scenario drastically, avoiding the formation of singularities. See also Feng et al. (2024).
@@
Dymnikova proposes a model that encodes the gravitational information of the spherically symmetric vacuum solution in such a way that near the central singularity—where the gravitational tension and tidal forces of the vacuum black hole become infinite—the energy density in the Dymnikova model (DyM) remains finite. This premise will be crucial for the model we examine in this work. This finite density in the DyM behaves as a positive cosmological constant near the origin, resulting in a regular black hole solution with a de Sitter core. It is also worth noting that the DyM has garnered significant attention in recent years for addressing various issues in physics Paul (2023); Konoplya and Zhidenko (2024); Estrada and Muniz (2023); Alencar et al. (2023); Estrada and Aros (2019); Bueno et al. (2024).
@@
A special case of Lovelock theories is Pure Lovelock theory (PL) Cai and Ohta (2006). As will be discussed below, this theory considers a single term in the Lagrangian. Pure Lovelock theory has drawn attention in recent years for various problems in physics Singha and Biswas (2024); Paithankar and Kolekar (2023); Shaymatov and Dadhich (2022). See also Chakraborty and Dadhich (2018, 2015) . As we will see below, the value of the Kretschmann scalar for the vacuum solution depends on the power $n$ of the curvature. This makes it interesting to study the behavior of a non-vacuum model that encodes the gravitational information of the empty geometry in an analogous way to the Dymnikova model.
Pure Lovelock is a theory characterized by specific properties that distinguish it from general Lovelock theory and all other higher derivative theories: vacuum solutions of the motion equations in Pure Lovelock theory are doubly degenerate for even $n$, a situation that does not occur in General Relativity where $n=1$. This also raises an intriguing question about what these degenerate solutions represent in the non-vacuum case and whether they allow for transverse geometries that differ from spherical symmetry. Additionally, it is well known that General Relativity has a non-trivial vacuum solutions, where the gravitational potential of the form $-\bar{M}/r^{d-2n-1}$ does not depend radially on its denominator when $d=3$ (i.e., $d=2n+1$ where $n=1$). An interesting feature is that pure Lovelock theory retains this property for $d=2n+1$ with $n>1$. Related to this, references Dadhich (2016); Dadhich et al. (2012); Camanho and Dadhich (2016) discuss a universal property of pure Lovelock theory, highlighting its kinematic nature in critical odd dimensions $d=2N+1$. This property raises the question of what the physical interpretation would be in the non-vacuum case, i.e., when $d=2n+1$ in the presence of high curvature terms. Another remarkable property of Lovelock theory is the existence of bound orbits in higher dimensions Dadhich et al. (2013).
@@ IV Our regular black hole solution in Pure Lovelock gravity
    $\lim_{r\gg a}\mu(r)_{\gamma=1}=1-\left(\frac{2(n-1)!M}{r^{d-2n-1}}\right)^{\frac{1}{n}},$ (28)
    $\lim_{r\gg a}\mu(r)_{\gamma=-1}=-1+\left(\frac{2(n-1)!M}{r^{d-2n-1}}\right)^{\frac{1}{n}}$ (29)
@@
    $\lambda(r)=\mu(r)_{\gamma=1}\Big{|}_{r\ll a}1-\left(\frac{2M}{a^{d-1}\cdot n}\right)^{\frac{1}{n}}r^{2}.$ (30)
@@
  • Also we can check that the solution with $\gamma=-1$ and even $n$ behaves near the origin as
    $\mu(r)_{\gamma=-1}\Big{|}_{r\ll a}-1+\left(\frac{2M}{a^{d-1}\cdot n}\right)^{\frac{1}{n}}r^{2}.$ (31)
    Therefore, this later solution possesses an Anti de Sitter core, representing a regular black hole.
@@ V.1.1 $d=2n+1$
Figure 1: Left side: Parameter of $M$ for $a=1$. Right side: Function $\mu$ for $a=1$, $M=0.6$.
@@ V.1.2 $d>2n+1$
Figure 2: Parameter of $M$ for $a=1$ and odd $n$.
Figure 3: Parameter of $M$ for $a=1$ and even $n$.
@@
Figure 5: Left side: Function $\mu$ for $a=1$, $n=2$, $d=7$, $M(r_{*})=1.28$. Right side: Function $\mu$ for $a=1$, $n=3$, $d=10$, $M(r_{*})=1.28$.
@@ V.2.2 $d>2n+1$
Figure 7: Left side: Function $\mu$ for $a=1$, $n=2$, $d=7$, $M(r_{*})=1.29$. Right side: Function $\mu$ for $a=1$, $n=4$, $d=10$, $M(r_{*})=0.215$.
@@
V.3 A brief discussion about the topology
Given that in our solutions the mass function is such that the spacetime behaves differently both near and far from the origin, it is of physical interest to test the topology of the spacetime at different locations. In this regard, initially, in reference Borde (1997), using the analogy with the Reissner-Nordström spacetime, conditions for the so-called Borde’s theorem were established, which leads to the fact that regular black holes with spherical symmetry (and consequently compact on the event horizon) in their cross-section are associated with a change in topology. Subsequently, in the references Bargueño (2020); Melgarejo et al. (2020); Bargueño (2021) the topology of four-dimensional regular black holes with (A)dS cores and other types of cores and with spherical symmetry in their cross-section was studied. In reference Bargueño (2020), it was found that the geometry of spacelike slices of spherically symmetric and (locally) static regular solutions is $S^{3}$ for a de Sitter core and $H^{3}$ for an AdS core. In that reference, it was pointed out that the regular black holes with dS core satisfying Borde’s theorem Borde (1997) has the topology of $S^{3}$ in their cores, assuming the compact slice is smooth and simply connected. On the other hand, there is a transition between the topology of the core and the topology far from it, from $S^{3}$ to ×$\mathcal{R}\times S^{2}$. Related to the latter, in reference Melgarejo et al. (2020), a way of representing the spacetime of a regular black hole was addressed, providing an identification of the regions corresponding to the origin in the usual Penrose diagram of Reissner-Nordström. This allows the spacetime to be described as causally simple towards the future and geodesically complete towards the future. It is also shown that all the requirements of Borde’s theorem are satisfied by imposing the null energy condition. Therefore, there must be some compact slice in the causal future of the eventually trapped surface. This spacelike identification implies that the topology of the slices in the inner region of the regular black hole is $S^{3}$, where the metric $ds^{2}=r_{-}d\chi^{2}+r^{2}(\chi)d\Omega^{2}$ is such that $r(\chi)$ fluctuates between 00 and $r_{-}$. This type of identification is also consistent with the aforementioned transition in topology.
In this subsection, we will have a brief discussion on the topology associated with the cases described in the previous subsections.
V.3.1 Branch with $\gamma=1$ for both even and odd $n$ for $d>2n+1$
As mentioned before, the case where $d=2n+1$ does not represent a black hole; therefore, the case of interest corresponds to $d>2n+1$. As we can observe in equation (30), the core corresponds to a dS space. In order to test the topology, we will follow a strategy analogous to that in reference Bargueño (2020). Since we are interested in testing whether the topology of the core corresponds to a sphere $S^{d-1}$, we will assume that the cross section of a coordinate system in any spacelike slice is constant and positively normalized to unity, $\gamma=1$. However, a more detailed study, such as the one in reference Bargueño (2020), may be required, which goes beyond the scope of this work. Such a coordinate system can be written as
$ds^{2}=\frac{dr^{2}}{\lambda(r)}+r^{2}d\Omega_{d-2}.$ (35)
where $\lambda(r)$ is given by (30). This last equation can be rewritten as:
$\lambda(r)=1-\frac{r^{2}}{l_{eff}}=1-\frac{R_{dS}}{d(d-1)}r^{2}=1-\frac{2}{(d-2)}\frac{\Lambda_{eff}}{(d-1)}r^{2}$ (36)
where $l_{eff}$, $R_{dS}$, and $\Lambda_{eff}$ represent the effective radius, Ricci scalar, and effective cosmological constant for the de Sitter spacetime, respectively. Intuitively motivated by reference Bargueño (2020), the following coordinate transformation is proposed:
$1-\frac{R_{dS}}{d(d-1)}r^{2}=\cos^{2}(\chi)\Rightarrow r=\left(\frac{d(d-1)}{R_{dS}}\right)^{1/2}\sin(\chi)$ (37)
with this change of coordinate, the line element (35) is:
$ds^{2}=\frac{d(d-1)}{R_{dS}}\left(d\chi^{2}+\sin^{2}(\chi)d\Omega_{d-2}\right)$ (38)
And therefore, the topology of the spacelike slices is described by either $S^{d-1}$ (corresponding to a de Sitter core with $R_{dS}>0$). On the other hand, we can check in equation (28) and figure 4 that, given the form of our mass function, the space time far from the origin behaves as the vacuum solution whose topology corresponds to ×$\mathcal{R}\times S^{d-2}$. Thus, we can infer that in our case, there is a transition in topology from an $S^{d-1}$ near the core to a topology of ×$\mathcal{R}\times S^{d-2}$ far from it. This result in $4D$ reduces to that in reference Bargueño (2020)
V.3.2 Branch with $\gamma=-1$ for even $n$
First of all, as pointed out previously, the case $d=2n+1$ does not have an inner horizon associated with a potentially unstable core. On the other hand, for $d>2n+1$, there is a black hole with both an event horizon and a cosmological horizon. As we can observe in equation (31), the solution has an AdS core. This feature makes it intriguing to test whether or not there is a topological transition analogous to the previous case, where now the cross-section corresponds to a hyperboloid spacetime. However, the fact that in our case the cross-section corresponds to a compact hyperboloid, together with the existence of a cosmological horizon, implies that the analogous identification with the Reissner-Nordström spacetime is not direct and requires a deep study, which is beyond the scope of this work and could be addressed in future studies.
@@ VI.1.1 $d>2n+1$
Figure 8: Left side: Temperature for $n=2$, $d=6$, $a=1$. Right side: Temperature for $n=3$, $d=8$, $a=1$.
@@ VI.2.2 $d>2n+1$
Figure 9: Left side: Temperature for $n=2$, $d=5$, $a=1$. Right side: Temperature for $n=4$, $d=9$, $a=1$.
@@ VIII.2.1 $d=2n+1$
Figure 10: Left side: Specific heat for $n=2$, $d=6$, $a=1$. Right side: Specific heat for $n=4$, $d=9$, $a=1$
@@ IX Discussion and summarize
Moreover, it is worth mentioning that, following references Borde (1997); Bargueño (2020); Melgarejo et al. (2020); Bargueño (2021), we have established that for the branch with a spherically symmetric cross-section for $d>2n+1$, near the de Sitter core, the topology of the spacelike slices is described by an $S^{d-1}$ sphere (corresponding to a de Sitter core with $R_{dS}>0$). On the other hand, given the form of the mass function, the spacetime far from the origin behaves as the vacuum solution, whose topology corresponds to ×$\mathcal{R}\times S^{d-2}$. Thus, we can infer that in this case, there is a transition in topology from an $S^{d-1}$ near the core to a topology of ×$\mathcal{R}\times S^{d-2}$ far from it. This result in $4D$ reduces to that in reference Bargueño (2020).